OpenStudy (anonymous):

Help Please!!!! (: For x in [-13, 15] the function f is defined by f(x)=x^6(x-6)^7 1.) On which two intervals is the function increasing. 2.) Find the region in which the function is positive. 3.) Where does the function achieve it's minimum?

OpenStudy (loser66):

first off, take derivative of f(x) what do you get? second, consider the sign of the f'(x) third, find the critical points, plug them back to f(x) , which smallest is minimum

OpenStudy (anonymous):

\[f \prime (x)= 7(x-6)^6*(x^6)+6(x-6)^7*(x^5)\]

OpenStudy (anonymous):

I solved for x and got x=0 x=36/16 and x=6

OpenStudy (loser66):

ok, then ? consider the sign of f'(x)

OpenStudy (anonymous):

what about the sign f'(x)?

OpenStudy (loser66):

hey, x =0, x =6 and x = 36/13, not 36/16

OpenStudy (loser66):

wait few second, I don't know whether it makes sense to you or not, I just post. I will scan it for you

OpenStudy (anonymous):

I see where I went wrong with the 36/16 it's 36/13. haha and ok. (:

OpenStudy (anonymous):

just to butt in for a second, it might be easier to analyze the sign of the derivative if you write it in factored form as \( x^5(x-6)^6 (13 x-36)\)

OpenStudy (anonymous):

Wow! okay that was an easy way to do the problem by making that table! Thank you so much! I really appreciate the help!

OpenStudy (loser66):

that means you understand my way? ahahaaa... good

OpenStudy (anonymous):

yes. sorta took me a while to understand the table. haha.

OpenStudy (anonymous):

Could you help me with another similar problem. I understand the process. but the derivative has no solutions.

OpenStudy (loser66):

don't promise, just post

OpenStudy (loser66):

if I can, i will